OpenStudy (anonymous):

What is the maximum number of electrons in the following energy level n=4? A. 2 B. 18 c. 8 D. 32

OpenStudy (stmoney):

4s2, 3d10, 4p6. 2 + 10 + 6 = 18

OpenStudy (ipwnbunnies):

3d10 isn't part of the 4th principal energy level...

OpenStudy (stmoney):

then 2 + 10 + 6 + 14 = 32

OpenStudy (ipwnbunnies):

4s2 + 4p6 + 4d10 + 4f14. Energy level n=4 can hold up to 32 electrons.

OpenStudy (anonymous):

:o